# On the Metallicity Gradients in the Galactic Disk using Open Clusters

### v1

@@ Abstract
We study the metallicity distribution and evolution in the Galactic disk based on the largest sample of open star clusters in the Galaxy. From the catalogue of 1879 open clusters in the range of Galactocentric distance () from 4 to 20 kpc, we investigate the variation of metallicity in the Galactic disk as functions of , vertical distance (), and ages of the clusters. In the direction perpendicular to the Galactic plane, variation in metallicity is found to follow a stepped linear relation. We estimate a vertical metallicity gradient of dex kpc for kpc, and dex kpc for kpc. On average, metallicity variations above and below the Galactic plane are found to change at similar rates. The change in metallicity in the radial direction is also found to follow a two-function linear relation. We obtain a radial metallicity gradient of dex kpc for kpc, and dex kpc for kpc which clearly shows a strong variation in the metallicity gradient when moving from the inner to the outer Galactic disk. Age-metallicity relation (AMR) is found to follow a steeper negative slope of dex Gyr for clusters older than 240 Myr, however, there is some hint of positive metallicity age gradient for younger clusters.
@@ 5 Age-Metallicity Relation
where, C is the point of intersection, b and b are the [Fe/H]-axis intercepts, and m and m are slopes for the two functions. The coefficients of the fitted regressions (along with the point of intersection C and corresponding standard errors) were determined using iterative least square estimation by treating b, m, C, and m as variables while assuming b = (m). Based on the distribution in Figure 3, we assumed the initial values of b, m, C, and m as 0, 0, 8.5, and -0.02, respectively. Estimated values of coefficients from each iteration were adopted as inputs for the next iteration. Along with minimizing the mean error, we also repeated the iterations until the difference between the estimated coefficients and the corresponding adopted coefficients was less than . In addition, for each of the iterations, we also checked the distribution of fitted metallicity residual (i.e., observed metallicity value minus predicted model value) as a function of and found that for the final iteration, the slope and the intercept to this distribution were less than . The final fitted function to our OCs data is shown as the red line in Figure 3. The coefficients of the fitted functions are also provided in the legend of Figure 3 in the form (b, m, C and m). Based on the sample OC, the two linear functions intersect at (i.e., an age of about 240 Myr), and age-metallicity gradients are given as:
@@ 6.1 Vertical metallicity gradient
where the negative slope indicates that metallicity decreases as we move away from the Galactic mid-plane. These average vertical-metallicity gradients over a large distance are in agreement with the previous studies. For example, Chen et al. [2003] found a vertical metallicity gradient of -0.295 0.050 dex kpc using a sample of 118 OCs. Through a sample of 40,000 stars with low-resolution spectroscopy over 144 lines of sight, Schlesinger et al. [2014] found a vertical metallicity gradient of dex kpc in different [/Fe] subsamples. However, as readily evident from both panels in Figure 6, a single linear fit is insufficient to explain the full trend in metallicity as a function of vertical distance. For kpc, the metallicity decreases rapidly while at the larger height, the change is relatively small.
To get a more accurate estimate for the vertical-metallicity gradient and also to find the vertical distance at which the radial-metallicity gradient changes significantly, we fitted the data with a combination of two linear regressions and the coefficients of the fitted functions are determined using iterative least square estimation following the procedure used in Section 5. Based on the distribution in Figure 6, we assumed the initial values of b1, m1, C, and m2 as 1.0, 0.4, 1.0, and 0.0, respectively. The final fitted function is shown as the red line in Figure 7. The coefficients of the fitted functions are also provided in the legends of figure 7 in the form (b, m, C, m), where b and m are the y-axis intercept and slope for the first linear function, m is the slope for the second linear function, and C is the point where these two function intersects. From the least square fitting, it is found that the two linear functions intersect at kpc and vertical-metallicity gradients are described as follows:
@@ 6.2 Radial metallicity gradient
The radial metallicity gradient is another important information to understand the chemical evolution of the Galactic disk and in turn the evolution of the Galaxy as the distribution of metallicity is not homogeneous across the Galaxy. It has been found that the metallicity in cluster population shows a decreasing trend with increasing distance from the GC (Wu et al. [2009]; Pancino et al. [2010a]; Yong et al. [2012b]; Donati et al. [2015b]; Magrini et al. [2015]; Carrera et al. [2019]; Donor et al. [2020]; Zhong et al. [2020]; Zhang et al. [2021]; Myers et al. [2022a]; Spina et al. [2022]; Netopil et al. [2022]; Magrini et al. [2023]). The radial metallicity gradient and its evolution with age are among the most critical empirical constraints that one can put on the Galactic chemical evolution models. Most of these models show that the formation of clusters strongly influences the appearance and development of radial metallicity gradients (Chiappini et al. [2001]) and the precise value of the metallicity gradient in the Galactic disk is an important parameter to constrain the chemical evolution models. The existence of such gradient across the Milky Way disk is well established through the observations of HII regions, disk stars, hot stars, star clusters, planetary nebula, Cepheid variables, and field stars (Chen et al. [2003]; Maciel et al. [2010]) and OCs (Carrera et al. [2019]; Donor et al. [2020]; Zhong et al. [2020]; Zhang et al. [2021]; Myers et al. [2022a]; Spina et al. [2022]; Netopil et al. [2022]; Magrini et al. [2023]). An average gradient of about dex kpc is observed in the Milky Way disk for most of the elements, e.g., O, S, Ne, Ar, and Fe. This magnitude of the observed gradients constrains the various parameters in the chemical evolution model, such as the time scales of star formation and infall (Prantzos and Aubert [1995]) or any variations of the stellar initial mass function properties with metallicities (Chiappini et al. [2001]).
@@
To find radial-metallicity gradients and the radial distance at which the radial-metallicity gradient changes, we fitted the data with a combination of two linear functions, and the coefficients of the fitted functions are determined using iterative least square estimation adopting the procedure followed in Section 5. Based on the distribution in Figure 8, we assumed the initial values of b1, m1, C, and m2 as 1.0, -0.05, 12.0, and -0.03, respectively. The final fitted function is shown as the red line in the figure. The coefficients of the fitted functions are also provided in the legends of the figure in the form (b, m, C, m), where b and m are the y-axis intercept and slope for the first linear function, m is the slope for the second linear function, and C is the point where these two function intersects. The two linear functions intersect at of kpc and gradients in - metallicity distributions are found as:
@@
Our estimated radial metallicity gradients are in close agreement with some of the recent determinations of metallicity gradients derived using samples of OCs. A comparison of radial metallicity gradients from some of the recent studies based on OCs along with our estimates is provided in Table 4. Most of these studies suggest a radial metallicity gradient of about -0.06 dex kpc. The radial metallicity gradient provides vital information on radial migration which plays an important role in the redistribution of stellar populations, particularly the older ones, in our Galaxy. It is believed that radial migration in OCs may be the reason for the flattening of the radial metallicity gradient over a period of time (Zhang et al. [2021]; Viscasillas Vázquez et al. [2022]). It is believed that there is a deficiency of low-metallicity clusters in the inner disk migrating from the outer disk, as the chance of survival in the high Galactic potential of the inner disk is low. On the other hand, clusters from the more metal-rich inner Galactic disk can migrate farther into the outer disk where the potentials of the spiral arm and bar are weaker, resulting in the enhancement of the mean metallicity of the outer disk. As a consequence, the radial metallicity gradient is steeper in the inner disk while flattening out towards the large galactocentric distance. Various earlier studies using different tracers such as planetary nebulae, classical Cepheids, globular clusters etc., also suggested that the radial metallicity gradient becomes slightly flatter with time ([e.g. Friel et al., 2002; Chen et al., 2003; Maciel and Costa, 2009; Luck et al., 2011; Genovali et al., 2014; da Silva et al., 2023]). It was contemplated by Toyouchi and Chiba [2014] that the radial metallicity gradient was positive at the time of formation of the thick disk which subsequently became negative during the transition phase of disk formation from thick to thin disk. It became flatter by the time of the formation of the thin disk. They credited this evolution of disk to the gas in-fall history having a shorter time scale in the inner disk and a relatively longer time scale in the outer disk, which is often called the ’inside-out’ scenario in the disk formation.
@@ 6.4 Vertical evolution of radial metallicity gradient
The study of the relation between the metallicity and the location of the cluster on the Galactic disk is an important tool for the study of the structure formation and evolution of the Galaxy ([e.g., Zhong et al., 2020]). We also investigated the evolution of radial metallicity gradients in the vertical direction of the Galactic plane as clusters are widely distributed in the vertical direction of the Galactic disk. The effect of scale height on the rate of change in metallicity variation with the has been analyzed by plotting the slope of the radial metallicity gradients as a function of the absolute value of which is shown in Figure 9. for our sample ranges from 0 to about 1.8 kpc (excluding one lone cluster located at about 2.6 kpc). While most of the clusters are located near the Galactic plane, there are fewer clusters at larger vertical distances. Therefore, we considered a varying bin size in the scale. We considered a bin width of 50 pc for , 100 pc for , and 500 pc for by making sure that there are enough clusters in the selected bins to get a proper estimate of radial metallicity gradient. As shown in Figure 9, the variation of as a function of absolute vertical distance for our sample of OCs follows an increasing trend with an increase in . Estimated standard errors in and mean are shown as black coloured cross bars. At larger vertical distances, the estimated errors are larger because of the smaller sample size in corresponding vertical distance bins. A linear fit to the distribution is shown as a thick red line and corresponding coefficients are shown in the legends. The radial metallicity gradient () as a function of vertical distance from the Galactic plan is found to vary at a rate of dex kpc kpc, suggesting a higher rate of change of metallicity with at larger vertical distance from the Galactic plane. The radial metallicity gradient at the centre of the Galatic plane is estimated as dex kpc which is in agreement with our previous estimate in Section 6.2.
@@ 7 Discussion and Conclusions
Using our sample of clusters, we further explored the vertical and radial metallicity gradients in the Galactic disk. Metallicity was found to follow a stepped variation with vertical distance from the Galactic plane. Near the Galactic plane, with kpc, we estimated the vertical metallicity gradient of dex kpc, while for large vertical distance having kpc, we found a lower vertical metallicity gradient of dex kpc. The lower metallicity gradient at large vertical distances compared to one at smaller vertical distances agrees with the Galactic chemical evolution models. We found that most of the OCs at large vertical distances are older compared to the majority of the clusters located near the Galactic plane. This difference in the ages of clusters from the two vertical regions is believed to be the main reason for the flatter vertical metallicity gradient at large vertical distances compared to the steep vertical metallicity gradient at smaller vertical distances.
Similar to the vertical direction, the change in metallicity in the radial direction is also found to follow a stepped linear relation. For a radial distance between about 4.0 to 12.8 kpc, we found a radial metallicity gradient () of dex kpc while for radial distance between about 12.8 to 20.5 kpc, we found a much smaller radial metallicity gradient of dex kpc. Thus the OCs in the outer Galactic disc are generally more metal-poor than the OCs in the inner Galactic disc and in the solar neighborhood. Although a shallower metallicity gradient in the region 12.8-20.5 kpc may be biased due to the relatively smaller number of OCs at larger Galactocentric distances but it could also be the result of radial migration of clusters in the Galactic disk (Zhang et al. [2021]). Using a smaller sample of 295 OCs within a Galactocentric distance of 7-15 kpc, Zhong et al. [2020] reported a steeper slope of -0.2520.039 dex kpc. It should also be noted that a significant variation in the slope as well as the turn-off point in the radial metallicity gradient among different studies, comes from the choice of the cluster sample, selected range of , and unequal vertical heights. Overall, our radial metallicity gradient estimates agree with most of the recent studies (Reddy et al. [2020]; Zhang et al. [2021]; Myers et al. [2022a]).
@@
We further studied the variation of radial metallicity gradient with distance from the Galactic plane and found that the radial metallicity gradient linearly increases with an increase in vertical distance and obtained a radial metallicity gradient slope of dex kpc kpc as a function of vertical distance from the Galactic plane. This agrees with the Galactic evolutionary models, for example, see Toyouchi and Chiba [2014] and references therein. In the case of a thin disk, which has a scale height of about 300 pc, the radial metallicity gradient is highly negative even though linearly increasing with vertical distance from the Galactic plane. However, for the thick disk (having a typical scale height of about 900 pc), the radial metallicity gradient is slightly high and approaches zero at about 1 kpc. Toyouchi and Chiba [2014] suggested that the radial metallicity gradient was positive at the time of formation of the thick disk but subsequently became negative during the transition phase of the disk formation from thick to thin disk. The gradient became flatter by the time of the formation of the thin disk. This change in radial metallicity gradient with vertical distance is believed to be related to the gas-infall history in the Galaxy. A large negative radial metallicity gradient near the Galactic plane (i.e., in the thin disk) but a higher gradient in the case of the thick disk (i.e., large vertical distance) can be explained by shorter and longer time scales, respectively, for the gas in-fall.

### v2

@@ Abstract
We study the metallicity distribution and evolution in the Galactic disk based on the largest sample of open star clusters in the Galaxy. From the catalogue of 1879 open clusters in the range of Galactocentric distance () from 4 to 20 kpc, we investigate the variation of metallicity in the Galactic disk as functions of , vertical distance (), and ages of the clusters. In the direction perpendicular to the Galactic plane, variation in metallicity is found to follow a stepped linear relation. We estimate a vertical metallicity gradient of dex kpc-1 for kpc, and dex kpc-1 for kpc. On average, metallicity variations above and below the Galactic plane are found to change at similar rates. The change in metallicity in the radial direction is also found to follow a two-function linear relation. We obtain a radial metallicity gradient of dex kpc-1 for kpc, and dex kpc-1 for kpc which clearly shows a strong variation in the metallicity gradient when moving from the inner to the outer Galactic disk. Age-metallicity relation (AMR) is found to follow a steeper negative slope of dex Gyr-1 for clusters older than 240 Myr, however, there is some hint of positive metallicity age gradient for younger clusters.
@@ 5 Age-Metallicity Relation
where, C is the point of intersection, b1 and b2 are the [Fe/H]-axis intercepts, and m1 and m2 are slopes for the two functions. The coefficients of the fitted regressions (along with the point of intersection C and corresponding standard errors) were determined using iterative least square estimation by treating b1, m1, C, and m2 as variables while assuming b2 = (m). Based on the distribution in Figure 3, we assumed the initial values of b1, m1, C, and m2 as 0, 0, 8.5, and -0.02, respectively. Estimated values of coefficients from each iteration were adopted as inputs for the next iteration. Along with minimizing the mean error, we also repeated the iterations until the difference between the estimated coefficients and the corresponding adopted coefficients was less than . In addition, for each of the iterations, we also checked the distribution of fitted metallicity residual (i.e., observed metallicity value minus predicted model value) as a function of and found that for the final iteration, the slope and the intercept to this distribution were less than . The final fitted function to our OCs data is shown as the red line in Figure 3. The coefficients of the fitted functions are also provided in the legend of Figure 3 in the form (b1, m1, C and m2). Based on the sample OC, the two linear functions intersect at (i.e., an age of about 240 Myr), and age-metallicity gradients are given as:
@@ 6.1 Vertical metallicity gradient
where the negative slope indicates that metallicity decreases as we move away from the Galactic mid-plane. These average vertical-metallicity gradients over a large distance are in agreement with the previous studies. For example, Chen et al. [2003] found a vertical metallicity gradient of -0.295 0.050 dex kpc-1 using a sample of 118 OCs. Through a sample of 40,000 stars with low-resolution spectroscopy over 144 lines of sight, Schlesinger et al. [2014] found a vertical metallicity gradient of dex kpc-1 in different [/Fe] subsamples. However, as readily evident from both panels in Figure 6, a single linear fit is insufficient to explain the full trend in metallicity as a function of vertical distance. For kpc, the metallicity decreases rapidly while at the larger height, the change is relatively small.
To get a more accurate estimate for the vertical-metallicity gradient and also to find the vertical distance at which the radial-metallicity gradient changes significantly, we fitted the data with a combination of two linear regressions and the coefficients of the fitted functions are determined using iterative least square estimation following the procedure used in Section 5. Based on the distribution in Figure 6, we assumed the initial values of b1, m1, C, and m2 as 1.0, 0.4, 1.0, and 0.0, respectively. The final fitted function is shown as the red line in Figure 7. The coefficients of the fitted functions are also provided in the legends of figure 7 in the form (b1, m1, C, m2), where b1 and m1 are the y-axis intercept and slope for the first linear function, m2 is the slope for the second linear function, and C is the point where these two function intersects. From the least square fitting, it is found that the two linear functions intersect at kpc and vertical-metallicity gradients are described as follows:
@@ 6.2 Radial metallicity gradient
The radial metallicity gradient is another important information to understand the chemical evolution of the Galactic disk and in turn the evolution of the Galaxy as the distribution of metallicity is not homogeneous across the Galaxy. It has been found that the metallicity in cluster population shows a decreasing trend with increasing distance from the GC (Wu et al. [2009]; Pancino et al. [2010a]; Yong et al. [2012b]; Donati et al. [2015b]; Magrini et al. [2015]; Carrera et al. [2019]; Donor et al. [2020]; Zhong et al. [2020]; Zhang et al. [2021]; Myers et al. [2022a]; Spina et al. [2022]; Netopil et al. [2022]; Magrini et al. [2023]). The radial metallicity gradient and its evolution with age are among the most critical empirical constraints that one can put on the Galactic chemical evolution models. Most of these models show that the formation of clusters strongly influences the appearance and development of radial metallicity gradients (Chiappini et al. [2001]) and the precise value of the metallicity gradient in the Galactic disk is an important parameter to constrain the chemical evolution models. The existence of such gradient across the Milky Way disk is well established through the observations of HII regions, disk stars, hot stars, star clusters, planetary nebula, Cepheid variables, and field stars (Chen et al. [2003]; Maciel et al. [2010]) and OCs (Carrera et al. [2019]; Donor et al. [2020]; Zhong et al. [2020]; Zhang et al. [2021]; Myers et al. [2022a]; Spina et al. [2022]; Netopil et al. [2022]; Magrini et al. [2023]). An average gradient of about dex kpc-1 is observed in the Milky Way disk for most of the elements, e.g., O, S, Ne, Ar, and Fe. This magnitude of the observed gradients constrains the various parameters in the chemical evolution model, such as the time scales of star formation and infall (Prantzos and Aubert [1995]) or any variations of the stellar initial mass function properties with metallicities (Chiappini et al. [2001]).
@@
To find radial-metallicity gradients and the radial distance at which the radial-metallicity gradient changes, we fitted the data with a combination of two linear functions, and the coefficients of the fitted functions are determined using iterative least square estimation adopting the procedure followed in Section 5. Based on the distribution in Figure 8, we assumed the initial values of b1, m1, C, and m2 as 1.0, -0.05, 12.0, and -0.03, respectively. The final fitted function is shown as the red line in the figure. The coefficients of the fitted functions are also provided in the legends of the figure in the form (b1, m1, C, m2), where b1 and m1 are the y-axis intercept and slope for the first linear function, m2 is the slope for the second linear function, and C is the point where these two function intersects. The two linear functions intersect at of kpc and gradients in - metallicity distributions are found as:
@@
Our estimated radial metallicity gradients are in close agreement with some of the recent determinations of metallicity gradients derived using samples of OCs. A comparison of radial metallicity gradients from some of the recent studies based on OCs along with our estimates is provided in Table 4. Most of these studies suggest a radial metallicity gradient of about -0.06 dex kpc-1. The radial metallicity gradient provides vital information on radial migration which plays an important role in the redistribution of stellar populations, particularly the older ones, in our Galaxy. It is believed that radial migration in OCs may be the reason for the flattening of the radial metallicity gradient over a period of time (Zhang et al. [2021]; Viscasillas Vázquez et al. [2022]). It is believed that there is a deficiency of low-metallicity clusters in the inner disk migrating from the outer disk, as the chance of survival in the high Galactic potential of the inner disk is low. On the other hand, clusters from the more metal-rich inner Galactic disk can migrate farther into the outer disk where the potentials of the spiral arm and bar are weaker, resulting in the enhancement of the mean metallicity of the outer disk. As a consequence, the radial metallicity gradient is steeper in the inner disk while flattening out towards the large galactocentric distance. Various earlier studies using different tracers such as planetary nebulae, classical Cepheids, globular clusters etc., also suggested that the radial metallicity gradient becomes slightly flatter with time ([e.g. Friel et al., 2002; Chen et al., 2003; Maciel and Costa, 2009; Luck et al., 2011; Genovali et al., 2014; da Silva et al., 2023]). It was contemplated by Toyouchi and Chiba [2014] that the radial metallicity gradient was positive at the time of formation of the thick disk which subsequently became negative during the transition phase of disk formation from thick to thin disk. It became flatter by the time of the formation of the thin disk. They credited this evolution of disk to the gas in-fall history having a shorter time scale in the inner disk and a relatively longer time scale in the outer disk, which is often called the ’inside-out’ scenario in the disk formation.
@@ 6.4 Vertical evolution of radial metallicity gradient
The study of the relation between the metallicity and the location of the cluster on the Galactic disk is an important tool for the study of the structure formation and evolution of the Galaxy ([e.g., Zhong et al., 2020]). We also investigated the evolution of radial metallicity gradients in the vertical direction of the Galactic plane as clusters are widely distributed in the vertical direction of the Galactic disk. The effect of scale height on the rate of change in metallicity variation with the has been analyzed by plotting the slope of the radial metallicity gradients as a function of the absolute value of which is shown in Figure 9. for our sample ranges from 0 to about 1.8 kpc (excluding one lone cluster located at about 2.6 kpc). While most of the clusters are located near the Galactic plane, there are fewer clusters at larger vertical distances. Therefore, we considered a varying bin size in the scale. We considered a bin width of 50 pc for , 100 pc for , and 500 pc for by making sure that there are enough clusters in the selected bins to get a proper estimate of radial metallicity gradient. As shown in Figure 9, the variation of as a function of absolute vertical distance for our sample of OCs follows an increasing trend with an increase in . Estimated standard errors in and mean are shown as black coloured cross bars. At larger vertical distances, the estimated errors are larger because of the smaller sample size in corresponding vertical distance bins. A linear fit to the distribution is shown as a thick red line and corresponding coefficients are shown in the legends. The radial metallicity gradient () as a function of vertical distance from the Galactic plan is found to vary at a rate of dex kpc-1 kpc-1, suggesting a higher rate of change of metallicity with at larger vertical distance from the Galactic plane. The radial metallicity gradient at the centre of the Galatic plane is estimated as dex kpc-1 which is in agreement with our previous estimate in Section 6.2.
@@ 7 Discussion and Conclusions
Using our sample of clusters, we further explored the vertical and radial metallicity gradients in the Galactic disk. Metallicity was found to follow a stepped variation with vertical distance from the Galactic plane. Near the Galactic plane, with kpc, we estimated the vertical metallicity gradient of dex kpc-1, while for large vertical distance having kpc, we found a lower vertical metallicity gradient of dex kpc-1. The lower metallicity gradient at large vertical distances compared to one at smaller vertical distances agrees with the Galactic chemical evolution models. We found that most of the OCs at large vertical distances are older compared to the majority of the clusters located near the Galactic plane. This difference in the ages of clusters from the two vertical regions is believed to be the main reason for the flatter vertical metallicity gradient at large vertical distances compared to the steep vertical metallicity gradient at smaller vertical distances.
Similar to the vertical direction, the change in metallicity in the radial direction is also found to follow a stepped linear relation. For a radial distance between about 4.0 to 12.8 kpc, we found a radial metallicity gradient () of dex kpc-1 while for radial distance between about 12.8 to 20.5 kpc, we found a much smaller radial metallicity gradient of dex kpc-1. Thus the OCs in the outer Galactic disc are generally more metal-poor than the OCs in the inner Galactic disc and in the solar neighborhood. Although a shallower metallicity gradient in the region 12.8-20.5 kpc may be biased due to the relatively smaller number of OCs at larger Galactocentric distances but it could also be the result of radial migration of clusters in the Galactic disk (Zhang et al. [2021]). Using a smaller sample of 295 OCs within a Galactocentric distance of 7-15 kpc, Zhong et al. [2020] reported a steeper slope of -0.2520.039 dex kpc-1. It should also be noted that a significant variation in the slope as well as the turn-off point in the radial metallicity gradient among different studies, comes from the choice of the cluster sample, selected range of , and unequal vertical heights. Overall, our radial metallicity gradient estimates agree with most of the recent studies (Reddy et al. [2020]; Zhang et al. [2021]; Myers et al. [2022a]).
@@
We further studied the variation of radial metallicity gradient with distance from the Galactic plane and found that the radial metallicity gradient linearly increases with an increase in vertical distance and obtained a radial metallicity gradient slope of dex kpc-1 kpc-1 as a function of vertical distance from the Galactic plane. This agrees with the Galactic evolutionary models, for example, see Toyouchi and Chiba [2014] and references therein. In the case of a thin disk, which has a scale height of about 300 pc, the radial metallicity gradient is highly negative even though linearly increasing with vertical distance from the Galactic plane. However, for the thick disk (having a typical scale height of about 900 pc), the radial metallicity gradient is slightly high and approaches zero at about 1 kpc. Toyouchi and Chiba [2014] suggested that the radial metallicity gradient was positive at the time of formation of the thick disk but subsequently became negative during the transition phase of the disk formation from thick to thin disk. The gradient became flatter by the time of the formation of the thin disk. This change in radial metallicity gradient with vertical distance is believed to be related to the gas-infall history in the Galaxy. A large negative radial metallicity gradient near the Galactic plane (i.e., in the thin disk) but a higher gradient in the case of the thick disk (i.e., large vertical distance) can be explained by shorter and longer time scales, respectively, for the gas in-fall.